Hydrogen cyanide, HCN, is a poisonous gas. The lethal dose is approximately 300. mg HCN per kilogram of air when inhaled. The density of air at26 C is 0.00118 g/cm3 Part A Calculate the amount of HCN that gives the lethal dose in a small laboratory room measuring 12.0 ftx 15.0 ftx9.70t Express your answer to three significant figures and include the appropriate units. Part B Consider the formation of HCN by the reaction of NaCN (sodium cyanide) with an acid such as H2S04 (sulfuric acid) 2NaCN(s)+H2S04(aq)- Na2S04(aq)+2HCN(g) What mass of NaCN gives the lethal dose in the room? Express your answer to three significant figures and include the appropriate units. **Part C HCN forms when synthetic fibers containing Orlon or Acrilan burn. Acrilan has an empirical formula of CH2CHCN, so HCN is 50 9% of the formula by mass A rug in the laboratory measures 12.0x 13.0 ft and contains 30.0 oz of Acrilan fibers per square yard of carpet. If the rug burns, what mass of HCN will be generated in the room? Assume that the yield of HCN from the fibers is 20.0% and that the carpet is 38.0 % consumed Express your answer to three significant figures and include the appropriate units
Answer #1

Q!

HCN lethal dose = 300 mg / kg of air

D = 0.00118 g /cm3; T = 26°C = 299 K

if

V = 12*15*9.7 ft3 = 1746 ft3

V = (1746 ft3) * 28.3168 L / ft3 = 49441.21 L = 49441.21*10^3 mL

mass of air = D*V = ( 49441.21*10^3 mL)*0.00118 g/mL = 58340.6278 g of air

kg of air = 58340.6278/!000 = 58.34 kg of air

so we need --> mass of HCN --> 300 mg / kg of air * 58.34 kg = 17502 mg of HCN will make this toxic

B)

mass of NaCH will do lethal those

mol of HCN = mass/MW = (17502 /1000)/(27.0253 ) = 0.6476 mol of HCN

then...

2 mol of HCN = 2 mol of NaCN

0.6476 mol o fHCN = 0.6476 mol o fNaCN

mass of NaCN = mol*MW = 0.6476*49.0072 = 31.7370 g of NaCN will give a letahl those